OpenStudy (friendzone):

Describe the steps to solve 5x=3x+4

OpenStudy (anonymous):

your going to start by subtracting 3x from each side. (3x-3x=0), (5x-3x=2x) now you are left with 2x=4.. then, you are going to divide 2 on each side.. (2x/2=0), (4?2=2).. you get your answer of x=2.. plug in 2 where each of your x's where and you can check your answer.. 5x=3x=+4...........5(2)=3(2)+4...........{ 10=6+4 }